If tha point P( 4,3) is rotated through an angle of 450 in anti-clockwise direction about the origin, then coordinates of P in new position are ? Explain how ?

Grade:12th Pass

4 Answers

Aman Bansal
12 years ago

Dear Aditi,

The new coordinates (X,Y) are given as:

x=X Cosa - Y Sina

y= XCosa +Y Sina;

a = angle of rotation,

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Aditi Bhatnagar
12 years ago

But here the point is rotated and not the axis....so can we use the same formula as applicable to the rotation of axis???

Jit Mitra
12 years ago

z= 4+3i

Let the new co-ordinate after rotation be z1 .

 

z1 = z*e^(i.pi/4) = (4+3i)*(cos(pi/4) + i.sin(pi/4)) = 1/sqrt(2)* (4+3i)(1+i) = 1/sqrt(2)* (1+7i)

 

New Point = ( 1/sqrt(2), 7/sqrt(2) )

vishal lakha
12 years ago

it will remain (4,3)..lolz